# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v2

@@ 4.5 Limitations
Although our results are significant in demonstrating a viable path forward to scale up speech BCIs, there remain a number of limitations to the present work. We focused here on two downstream tasks: speech detection and voice classification. Ultimately, we would like to expand this work to predict full transcripts from brain recordings (i.e. brain-to-text). This has been achieved with surgical data [40, 60] but not yet convincingly with non-invasive methods like MEG or EEG [28]. Speech detection has played an important role in the development of full brain-to-text in a surgical context [40] and we hope may play a similar role for non-invasive methods. Prior work has further used voice classification as a stand in for phoneme classification [23], and we have been able to improve on these results here. In future work, we would like to expand this to all English phonemes. Secondly, while we have been able to demonstrate the utility of a few pretext tasks here, we do not claim to have exhausted the full set of useful tasks. Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings. For example, this paper did not make use of spatial features. Another limitation is our emphasis on heard speech over other types of speech, such as attempted or imagined speech. We hypothesise that the same methods presented here will generalise to these other varieties of speech, though this has yet to be shown. But, perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training. In sharing this work now, we believe that the current proof of concept will be sufficiently impactful to the field as we continue to actively scale up the datasets that we can leverage.
@@ 5 Conclusion
Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties. Yet, this promise has not materialised because the field has been blocked by its inability to scale up data to leverage deep learning. Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants. A handful of studies have shown weak signals towards alleviating these issues. But until now, no one has developed a general solution. We provided a unified method that leverages unlabelled data using generic pretext tasks that shows that all of these problems can be solved. We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks. Our method unlocks the potential of the bitter lesson, providing a general method to exploit more computation by using more data. We implore the research community to employ the vast quantities of data and compute available to realise this potential. If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
Acknowledgments and Disclosure of Funding
We would like to thank Botos Csaba for many insightful discussions and creative ideas which helped shaped the direction of this work. Thanks also to Minqi Jiang for an encouraging conversation on unsupervised representation learning, Mats W.J. van Es for technical assistance with the OSL library, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558.
[…]
Appendix A Supervised Heard Speech Decoding
While this work is primarily about scaling speech decoding, we also set the first baselines for speech detection and voicing classification from MEG with connected auditory speech stimuli. Therefore, in this section we provide fully supervised results where our architecture is trained to completion using all available labelled data in the corresponding dataset (except the validation and test splits) and compare this to a linear classifier baseline. These results are not comparable with our self-supervised models because the entire architecture is trained end-to-end. In our SSL work with shallow fine-tuning, only the final classifier is trained with labels.
Experiment Balanced accuracy
Speech detection Voicing
Armeni (linear) $0.5222\pm 0.0015$ $0.5207\pm 0.0007$
Armeni (ours) $\mathbf{0.7253}\pm 0.0005$ $\mathbf{0.5282}\pm 0.0006$
Gwilliams (linear) $0.5006\pm 0.0001$ $0.5166\pm 0.0002$
Gwilliams (ours) $\mathbf{0.5946}\pm 0.0016$ $\mathbf{0.5211}\pm 0.0005$
Table 3: Supervised baselines. Uncertainty is standard error of the mean over three random seeds.
The results in Table 3 show that we can achieve strong speech detection results on Armeni et al. [2] and Gwilliams et al. [24] with our architecture. The difference between the two may be due to a combination of different scanners, numbers of sensors (208 vs 269 magnetometers), varied difficulty in the audio stimulus, and more within-subject data in Armeni et al. [2]. We also achieve good results on voicing classification for both datasets, easily outperforming the linear models.
One potential way to improve on both supervised and self-supervised approaches is by using a temporally aware model such as an LSTM or an attentional model such as a Transformer as the classifier. We do not investigate this within this paper as it is out of scope. Rather, we choose to use simple classifiers as the purpose of this paper is to demonstrate the potential of scaling.
Appendix B Subject Conditioning
To investigate subject conditioning approaches, we pre-train then fine-tune our models on a subset of Gwilliams et al. [24], holding out three subjects for evaluation. We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation. We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation. For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
Subject Subject voicing balanced accuracy
conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
Chance-level $0.5$- -- $0.5$- --
Table 4: Subject conditioning improves generalisation on seen subjects but not novel subjects. We quote test accuracy at the point of best validation accuracy (early stopping). Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
Table 4 reveals that in-distribution subject generalisation is helped by using subject conditioning methods. Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best. However, for novel subject generalisation, using a subject conditioning method makes little difference. Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects. Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
Appendix C Aggregating Unlabelled Datasets
To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets. As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS [47] and Cam-CAN [48, 53]. We investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
Balanced accuracy
Labels Pre-training data (hours) Speech detection Voicing
Armeni MOUS (160 hours) $0.5736\pm 0.0016$ $0.5220\pm 0.0013$
Cam-CAN (159 hours) $\mathbf{0.5911}\pm 0.0009$ $\mathbf{0.5241}\pm 0.0004$
Cam-CAN + MOUS (319 hours) $0.5872\pm 0.0012$ $0.5236\pm 0.0016$
Gwilliams MOUS (160 hours) $0.5060\pm 0.0006$ $0.5175\pm 0.0008$
Cam-CAN (159 hours) $\mathbf{0.5118}\pm 0.0011$ $\mathbf{0.5196}\pm 0.0010$
Cam-CAN + MOUS (319 hours) $0.5116\pm 0.0017$ $0.5187\pm 0.0013$
Table 5: Combining two unlabelled datasets does not improve over the best study. We quote test accuracy at the point of best validation accuracy (early stopping). The uncertainty is the standard error of the mean calculated over three seeds.
The results in Table 5 suggest that combining these two datasets is not strictly better than the best dataset on its own. Pre-training on Cam-CAN alone leads to the best performance in all instances. This suggests that Cam-CAN is a better pre-training dataset than MOUS—a surprising result given that MOUS and Armeni et al. [2] share the same scanning hardware and MOUS includes data from subjects listening to speech while Cam-CAN has no such language tasks. This could be due to Cam-CAN being a cleaner dataset as we found artefacts can significantly affect learning. Combining Cam-CAN with MOUS performs marginally worse or the same as Cam-CAN alone. For all but one instance they are within the standard error of the mean of each other. The combined datasets always outperform MOUS on its own.
In general, further investigation is necessary to determine how datasets should be best aggregated. Increasing the number of datasets could enable the network to model the variance between them and eventually improve over the best singular dataset. This remains an open question as we were only able to study up to two datasets in this paper.
Appendix D Task Labels Are Hard To Scale; Generic SSL Scales Best
Figure 5: Cross-task semi-supervised learning does not improve generalisation. We pre-train with increasing amounts of labelled data from Armeni et al. [2], using speech labels when evaluating voicing and voicing labels for speech. We also use a constant 38 hours of unlabelled data in pre-training. No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). SSL is when pre-training with only the unlabelled data. The shaded area shows the standard error of the mean across three seeds.
Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding. However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias? This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training. Figure 5 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks. There also does not appear to be any indication of scaling within the range of labelled data available to us.
This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy. The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space. Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks. This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech; nor does it mean more labelled data will not scale. However, the point remains salient. Labelled data is more difficult to scale given its scarcity and the generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
Appendix E Experiment Details
To generate the data in Table 1, we pre-train our representations with non-overlapping sample windows from all subjects and sessions 001 to 009 of Armeni et al. [2] for 200 epochs with a $0.8:0.1:0.1$ train/val/test split. We shallow fine-tune with labels for 30 epochs using all subjects, but only session 010, which was held-out during pre-training.
@@ Appendix E Experiment Details
For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting. We pre-train only with Cam-CAN [48, 53]. We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects. We fine-tune with all data in Armeni et al. [2] and Gwilliams et al. [24] with the same train/val/test ratios as before.
For the results shown in Figure 5, we use the same setup as in Figure 3. In addition to the unlabelled data, we add labelled data from Armeni et al. [2], increasing the number of sessions used in the sequence 1, 3, 6, 9, 12, and 27. For this experiment, we calculate the number of labelled data hours using an estimate of 1 hour per session per subject.
Finally, for the results in Table 4, we pre-train and deep fine-tune to saturation with subjects 01-07 of Gwilliams et al. [24] and test zero-shot generalisation using held-out subjects 08-10. All other experimental details are the same as the previous two experiments.
@@
In all experiments, we use three randomly selected seeds for each pre-training and corresponding fine-tuning run. For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate. Therefore, each speech detection label represents a 0.1 second period of time.
Appendix F Hyperparameters
We conduct a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture. In all experiments in this section, we measure downstream performance (using shallow fine-tuning) with best validation accuracy for speech detection unless explicitly specified, training and evaluating only on data from subject 001 of Armeni et al. [2] to minimise training time, over three random seeds. All uncertainties in this section are standard deviations and not standard error of the mean as in the main body.
Table 6 shows the results of a search over the proportion $\rho$ of sensors to apply the phase and amplitude pretext task transformations to. We only search up to $\rho=0.5$ as anything above this is equivalent to applying the inverse transformation to a proportion $1-\rho$ of the sensors. We find that for phase shift prediction, $\rho=0.5$ is optimal, while for amplitude scale prediction $\rho=0.2$ is best.
Task $\rho$ Balanced accuracy
Phase shift prediction 0.1 $0.5452\pm 0.0048$
0.2 $0.5566\pm 0.0014$
0.3 $0.5578\pm 0.0021$
0.4 $0.5530\pm 0.0014$
0.5 $\mathbf{0.5655}\pm 0.0026$
Amplitude scale prediction 0.1 $0.5440\pm 0.0012$
0.2 $\mathbf{0.5796}\pm 0.0050$
0.3 $0.5625\pm 0.0063$
0.4 $0.5684\pm 0.0016$
0.5 $0.5650\pm 0.0026$
Table 6: Best $p$ for phase and amplitude pretext tasks.
In our architecture, we define $d_{\mathrm{backbone}}$ to be the bottleneck dimension of the cortex encoder (i.e. the embedding dimension) as well as the size of $d_{\mathrm{shared}}$. Balestriero et al. [3], using the results from work by Bordes et al. [4], suggest that increasing the backbone output dimension in self-supervised learning leads to better downstream transfer. We investigate this for our architecture by varying the encoder dimension in Table 7. Our results are consistent with their findings. Balanced accuracy is highest at the largest encoder dimension we evaluated. Encoder dimensions larger than $2048$ led to unstable training and were not feasible given the compute resources available to us.
$d_{\mathrm{backbone}}$ Balanced accuracy
$128$ $0.5775\pm 0.0037$
$300$ $0.5715\pm 0.0012$
$512$ $0.5835\pm 0.0016$
$1024$ $0.5800\pm 0.0050$
$2048$ $\mathbf{0.5987}\pm 0.0024$
Table 7: Optimal encoder dimension.
In Table 8, we show the effect of varying the length $w$, i.e. the window size of each data sample. Increasing the window to five seconds led to unstable losses in pre-training so we did not evaluate above a window length of two seconds. This is likely to be because increasing the window size reduces the number of training samples when only non-overlapping samples are used. For speech detection, there appears to be a U-shaped distribution, where very short and very long windows lead to better downstream accuracy. For voicing classification, the window lengths we evaluated seem to make little difference. However, windows smaller than 0.1 seconds may lead to significantly lower accuracy as most phonemes will occur for at least this length of time in the neural response [23].
Window length (s) Balanced accuracy
Speech detection Voicing classification
$0.2$ $0.5919\pm 0.0014$ $0.5033\pm 0.0014$
$0.4$ $0.5804\pm 0.0036$ $\mathbf{0.5077}\pm 0.0027$
$0.5$ $0.5715\pm 0.0012$ $0.5055\pm 0.0023$
$1.0$ $0.5790\pm 0.0012$ $0.5062\pm 0.0020$
$2.0$ $\mathbf{0.6116}\pm 0.0019$ $0.5065\pm 0.0011$
Table 8: Ideal sample window length.
We also evaluated several different configurations of the SEANet architecture for the cortex encoder, converging on convolutional blocks with channel dimensions $(512,512,512,512)$.
@@ Appendix F Hyperparameters
While these ablations indicate a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up. Our final architecture hyperparameters achieve a balance between the best values from our hyperparameter search and stable training. These values are detailed in Table 9.
@@
Optimizer AdamW [33]
@@
Appendix G Compute Resources
@@
Appendix H Licences For Datasets And Code
The Armeni et al. [2] dataset is distributed under CC-BY-4.0 while the Gwilliams et al. [24] dataset is distributed under the CC0 1.0 Universal licence. The Schoffelen et al. [47] dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute. The licence for the Cam-CAN [48, 53] dataset is unknown. The SEANet code adapted from Défossez et al. [14] is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.
Appendix I Broader Impacts
Decoding speech from non-invasive brain recordings is likely to bring about significant positive societal impacts. Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g. stammering). As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants. Nevertheless, the maturity of this technology could also present potential negative societal impacts. For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals. Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent. Moreover, there are possible long horizon effects associated with speech decoding research. Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society. In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality. Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse. Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.

### v3

@@ 4.6 Limitations
Although our results are significant in demonstrating a viable path forward to scale up speech BCIs, there remain a number of limitations to the present work. We focused here on two downstream tasks: speech detection and voice classification. Ultimately, we would like to expand this work to predict full transcripts from brain recordings (i.e. brain-to-text). This has been achieved with surgical data (Moses et al., 2021; Willett et al., 2023) but not yet convincingly with non-invasive methods like MEG or EEG (Jo et al., 2024). Speech detection has played an important role in the development of full brain-to-text in a surgical context (Moses et al., 2021) and we hope may play a similar role for non-invasive methods. Prior work has further used voice classification as a stand in for phoneme classification (Gwilliams et al., 2022), and we have been able to improve on these results here. In future work, we would like to expand this to all English phonemes. Secondly, while we have been able to demonstrate the utility of a few pretext tasks, we do not claim to have exhausted the full set of useful tasks. Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings. For example, this paper did not make use of spatial features. Another limitation is our emphasis on heard speech over other types of speech, such as attempted or imagined speech. We hypothesise that the same methods presented here will generalise to these other varieties of speech, though this has yet to be shown. But, perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training. In sharing this work now, we believe that the current proof of concept will be sufficiently impactful to the field as we continue to actively scale up the datasets that we can leverage.
@@ 5 Conclusion
Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties. This promise has not yet materialised because the field has been blocked by its inability to scale up data to leverage deep learning. Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants. A handful of studies have shown weak signals towards alleviating these issues. But until now, no one has developed a general solution. We provided a unified method that leverages unlabelled recordings data-efficiently using generic pretext tasks that shows that all of these problems can be solved. We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks. Our method unlocks the potential of the bitter lesson, providing a general method to exploit more computation by using more data. We implore the research community to employ the vast quantities of data and compute available to realise this potential. If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
Ethics Statement
In this work, we use data from studies that involve human subjects (Armeni et al., 2022; Gwilliams et al., 2023; Shafto et al., 2014; Taylor et al., 2017; Schoffelen et al., 2019). These datasets are public, cited, and have their own ethical approvals. The documentation for these is available with the publications for the respective datasets.
While there are clear positive impacts, we acknowledge that insights from neural speech decoding research may not all be beneficial. Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g. stammering). As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants. Nevertheless, the maturity of this technology could also present potential negative societal impacts. For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals. Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent. Moreover, there are possible long horizon effects associated with speech decoding research. Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society. In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality. Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse. Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.
Reproducibility Statement
In the supplementary materials, we have provided an anonymised code repository with instructions for reproducing our main experiments. We also include details on experiment design and setup (Section 4.1 and Appendix A), hyperparameters (Appendix B), and compute (Appendix C). While we attempt to be exhaustive with these details, any information not found directly in the main body or appendices can be located in the supplementary materials.
Appendix A Experiment Details
@@
We pre-train with non-overlapping sample windows from all subjects and sessions. We adjust the amount of unlabelled data used from Cam-CAN by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects.
When fine-tuning with Armeni et al. (2022), we hold out session 010 from all subjects during training and validation, using this for evaluation. Similarly, when fine-tuning with Gwilliams et al. (2023), we hold out session 1 from subjects 23, 24, 25, 26, and 27, using these sessions for evaluation only. As there is limited within-subject data in the latter dataset, we did not hold out a session from all subjects as before. For our novel subject experiments, we hold out subjects 1, 2, and 3 entirely and use the data for these subjects during evaluation. In Gwilliams et al. (2023), we note that they use four different tasks for each subject and their order is randomized between subjects. Both sessions for each task are repeats of the task. This means that while the recording itself is unseen, in this dataset, it is possible that heldout sessions use tasks that may have been seen in the training set.
@@
In all experiments, we always fine-tune to completion (usually around 300 epochs), taking the test metric at the best validation loss (early stopping). We use three randomly selected seeds for each pre-training and corresponding fine-tuning run. For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate. Therefore, each speech detection label represents a 0.1 second period of time.
Appendix B Hyperparameters
@@
We conducted a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture. While these ablations indicated a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up. Our final architecture hyperparameters achieve a balance between the best values from our hyperparameter search and stable training. These values are detailed in Table 3.
@@
Optimizer AdamW (Loshchilov & Hutter, 2019)
@@
Appendix C Compute Resources
@@
Appendix D Licences For Datasets And Code
The Armeni et al. (2022) dataset is distributed under CC-BY-4.0 while the Gwilliams et al. (2023) dataset is distributed under the CC0 1.0 Universal licence. The Schoffelen et al. (2019) dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute. The licence for the Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) dataset is unknown. The SEANet code adapted from Défossez et al. (2022) is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.